An isosceles triangle has a vertical angle of 120 degree and a base 40cm.
What will be the length of it's altitude.

Answer by Alan3354(69443)

An isosceles triangle has a vertical angle of 120 degree and a base 40cm.
What will be the length of it's altitude.
----------
Each "base angle" is 30 degs.
h = altitude
tan(30) = h/(base/2) = h/20 = sqrt(3)/3
h = 20*sqrt(3)/3 cm
=========================
it's = it is